OpenStudy (anonymous):

Whats the inverse of 2x+3 /5

OpenStudy (danjs):

It is the line , but reflected over y=x

OpenStudy (danjs):

just switch the y and x in y=2x+3/5 and resolve for the new y... that will give you the reflected line

OpenStudy (anonymous):

\[Suppose\ y=\frac{ 2x+3 }{ 5}\\ interchange\ y \ and\ x, i.e.\\ x=\frac{ 2y+3 }{ 5 }\\solve\ for\ y\\2y+3=5x\\2y=5x-3\\y=\frac{ 5x-3 }{ 2}\\put\ y= inverse \ of \ y=f(x)^-1 \ [assume]\\f(x)^-1=\frac{ 5x-3 }{ 2 }\ is\ desired\ inverse\ of\ y\]

OpenStudy (anonymous):

thanks all!

OpenStudy (anonymous):

cheers!